If Jaden is walking at a speed of 2 meters per second, and Ben is walking at a speed of 2.5 meters per second, and they live 20 meters away from each other, how many seconds will it take for them to meet.
Answer by mananth(16949)

If Jaden is walking at a speed of 2 meters per second, and Ben is walking at a speed of 2.5 meters per second, and they live 20 meters away from each other, how many seconds will it take for them to meet.
..
jaden speed = 2m/s
ben speed 2.5m/s
distance = 20 meters
they are walking towards each other
so add the speeds = 2.5+2=4.5m/s
time = distance /speed
20/2.5=8 seconds
...
m.ananth@hotmail.ca